OpenStudy (anonymous):

find an equation in slope-intercept form for the line through (2, -8), perpendicular to -7x+4y=-46

OpenStudy (anonymous):

Line given: -7x + 4y = -46 4y = 7x- 46 y = 7x/4 - 46 Slope of this line is 7/4. Slope of line perpendicular to this one will be -4/7. New line equation will be: y = -4x/7 + b (you can find b given this line passes through (2,-8). -8 = - 4*2/7 + b => b = -8 + 8/7 = -48/7 So, equation of that perpendicular line is: y = -4x/7 - 48/7 7y = -4x - 48

OpenStudy (anonymous):

thank you gt